You are to run a conduit diagonally across a parking lot that is 200 ft long and 60 ft wide.How much conduit will you need to complete this run? Round your answer to tenths.
Answer by nerdybill(7384)

Apply Pythagorean theorem
.
Let x = length of conduit
then
x^2 = 200^2 + 60^2
x^2 = 40000 + 3600
x^2 = 43600
x = 208.8 feet